OpenStudy (anonymous):

Common logarithm of 4,050,000?

OpenStudy (anonymous):

6.6075?

OpenStudy (anonymous):

is that a question or statment?

OpenStudy (anonymous):

I need to know if my answer if correct... I put a question mark after both

ganeshie8 (ganeshie8):

thats right. common logarithm is base 10. natural is base e you can work out the approximate like this : log(4.05*10^6) = log(4.05) + log(10^6) = 6 + log(4.05) = 6.xxxx

OpenStudy (anonymous):

well is 10 to the 6.6075 power 4,050,000 ?

OpenStudy (anonymous):

Thanks! :)

ganeshie8 (ganeshie8):

yw.. ...